OpenStudy (anonymous):

Which statement describes how work and power are similar? (2 points) You must know force and distance to calculate both. You must know energy and force to calculate both. You must know velocity and time to calculate both. You must know time and energy to calculate both.

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

i feel like its c

OpenStudy (anonymous):

wat do u think?

OpenStudy (igreen):

Hmm...hold on.

OpenStudy (anonymous):

actually i think its d

OpenStudy (anonymous):

so it is d

OpenStudy (igreen):

Power formula: \(\sf Power = \dfrac{Work~done}{Time~taken}\)

OpenStudy (igreen):

Work formula: \(\sf Work = Force \cdot Distance\)

OpenStudy (igreen):

So actually another way to write the formula for power is: \(\sf Power = \dfrac{Force \cdot Distance}{Time~taken}\)

OpenStudy (igreen):

I don't think it's C or D.

OpenStudy (anonymous):

so its a

OpenStudy (igreen):

Yep, I think sol.

OpenStudy (igreen):

*so